The future value of $9,000 in 6 years if invested at 6.3% per annum, compounding annually is (including cents; don’t use $ sign or comma separators):
Answer #1

P= 9000 14 = 6 years i= 6.3 % Future value = P(1+i) =9000 (10063) = $ 12985 . Anscer = $ 12985 or 12985
Answer #2

I need to have $59,000 saved in 7 years so that I can buy a car.  How much do I need to invest (to the nearest dollar) today if the appropriate interest rate is 13.6% per annum compounded semi-annually?